hi i am paul dinesh
I would like to know
1)what is the formula to find mass of 2 equivalent ofk Na2co3 ?

Grade:

1 Answers

mohit sengar
13 years ago

 

  eqivalent weight = molecular weight / n-factor

the n-factor of salts is the total no of cations on salt. 

as Na2CO3 is a salt and it has two cations or it has +2 positive charge , it means the n-factor of Na2CO3  is 2 

so , its equivalent weight is  =  106/2 gm 

but acc to question u hav asked for 2 equivalent  , therefore  , mass  = 2 × 106⁄2  = 106 gm